OpenStudy (anonymous):

Which of the following is an example of the Zero Product Property? A. If x⋅y=0 , then x=0 or y=0 B. If x+y=0 , then x=0 or y=0 C. Either x/ 0 =0 is correct or it can't be done without the Earth exploding. D. If zero is an even number then it is positive, otherwise zero is an odd number and is negative.

OpenStudy (cwrw238):

a product indicates a multiplication

OpenStudy (cwrw238):

which one means x multiplied by y?

OpenStudy (anonymous):

This could help - http://www.mathsisfun.com/algebra/zero-product-property.html

OpenStudy (anonymous):

option A...is correct.. option B invovles additive propeerty otion C is undefined option D is not much relavent to the statement..

OpenStudy (anonymous):

Example: if (x-5)(x-3) = 0 then (x-5) = 0 or (x-3) = 0

OpenStudy (anonymous):

Do you understand it now?